# Científicas (CSIC) and Universitat de València (UV), 46980 Valencia, Spain

### v1

@@ B.2 SILH Basis
We report here the projected sensitivity to universal operators in the SILH basis as well as a select number of non-universal operators which are defined differently to the Warsaw basis. Concretely, in addition to the operators introduced in Eqs. 16-19 and 24 -26, we focus on
[…]
@@
[…]
@@
$\mathcal{S}_{T}$ 80.62 $m_{W}$ 5.30 $\sigma_{WW}^{240}$ 80.62
$\mathcal{S}_{B+W}$ 26.57 $A^{Z}_{e}$ 3.59 $\sigma_{Zh}^{365}$ 26.57
$\mathcal{S}_{qD}^{(3)}$ 21.50 $R^{Z}_{\mu}$ 22.96 $R_{b}^{240}$ 26.24
$\mathcal{S}_{2W}$ 21.88 $A^{Z}_{e}$ 21.12 $R_{\tau}^{240}$ 25.37
$\mathcal{S}_{2B}$ 23.25 $A^{Z}_{e}$ 11.22 $R_{e}^{365}$ 23.55
$\mathcal{S}_{qD}^{(1)}$ 16.19 $A^{Z}_{e}$ 14.48 $R_{t}^{365}$ 18.28
$\mathcal{S}_{tD}$ 14.49 $A^{Z}_{e}$ 13.54 $R_{t}^{365}$ 16.64
$\mathcal{S}_{\gamma W}$ 9.29 $A^{Z}_{e}$ 7.76 $\sigma_{Zh}^{365}$ 10.22
$\mathcal{S}_{\gamma B}$ 9.16 $A^{Z}_{e}$ 3.95 $\sigma_{Zh}^{240}$ 9.23
$\mathcal{S}_{3W}$ 8.63 $A^{Z}_{e}$ 4.61 $\sigma_{WW}^{240}$ 8.79
$\mathcal{S}_{HW}$ 5.27 $m_{W}$ 4.46 $\sigma_{Zh}^{365}$ 5.81
$\mathcal{S}_{HB}$ 5.28 $m_{W}$ 3.45 $\sigma_{WW}^{240}$ 5.49
$\mathcal{S}_{H}$ 4.84 $m_{W}$ 2.33 $\sigma_{Zh}^{240}$ 4.89
$\mathcal{S}_{2G}$ 3.87 $m_{W}$ 3.02 $R_{e}^{365}$ 4.57
$\mathcal{S}_{B-W}$ 3.38 $A^{Z}_{e}$ 2.75 $\sigma_{Zh}^{365}$ 3.68
$\mathcal{S}_{g}$ 0.47 $m_{W}$ 2.19 $\mu_{b\bar{b}}$ 2.19
$\mathcal{S}_{6}$ 0.39 $m_{W}$ 0.83 $\sigma_{Zh}^{240}$ 0.82
@@
[…]

### v2

@@ B.2 SILH Basis
We report here the projected sensitivity to universal operators in the SILH basis as well as a select number of non-universal operators which are defined differently to the Warsaw basis. Concretely, in addition to the operators introduced in Eqs. 17-20 and 25 -27, we focus on
$\begin{array}[]{ccccc}\mathcal{O}_{\gamma W}=Q_{HW},&\hskip 14.22636pt\mathcal{O}_{\gamma B}=Q_{HB},&\hskip 14.22636pt\mathcal{O}_{3G}=Q_{3G},&\hskip 14.22636pt\mathcal{O}_{H}=-\frac{1}{2}Q_{H\square},&\hskip 14.22636pt\mathcal{O}_{6}=Q_{H},\end{array}$ (43)
@@
Table 5: Single Operator $2\sigma$ confidence intervals in the basis exceding $1\text{\,}\mathrm{TeV}$ from On- and Off-Pole observables, as well as their combined projected sensitivity. The Wilson Coefficients are renormalized at $\Lambda=$1\text{\,}\mathrm{TeV}$\$. We also report the most constraining measurement for each coefficient for On- and Off-Pole Observables.
@@
$\mathcal{S}_{T}$ 80.62 $m_{W}$ 6.41 $\sigma_{WW}^{240}$ 80.62
$\mathcal{S}_{2W}$ 21.88 $A^{Z}_{e}$ 25.01 $R_{\tau}^{240}$ 28.06
$\mathcal{S}_{qD}^{(3)}$ 13.02 $R^{Z}_{\mu}$ 26.52 $R_{b}^{240}$ 26.90
$\mathcal{S}_{B+W}$ 26.57 $A^{Z}_{e}$ 4.25 $\sigma_{Zh}^{365}$ 26.58
$\mathcal{S}_{2B}$ 23.25 $A^{Z}_{e}$ 13.13 $R_{e}^{365}$ 23.82
$\mathcal{S}_{qD}^{(1)}$ 16.18 $A^{Z}_{e}$ 16.82 $R_{t}^{365}$ 19.63
$\mathcal{S}_{tD}$ 14.49 $A^{Z}_{e}$ 15.73 $R_{t}^{365}$ 18.01
$\mathcal{S}_{\gamma W}$ 9.29 $A^{Z}_{e}$ 9.21 $\sigma_{Zh}^{240}$ 11.00
$\mathcal{S}_{\gamma B}$ 9.16 $A^{Z}_{e}$ 4.71 $\sigma_{Zh}^{240}$ 9.31
$\mathcal{S}_{3W}$ 8.63 $A^{Z}_{e}$ 5.52 $\sigma_{WW}^{240}$ 8.97
$\mathcal{S}_{HW}$ 5.27 $m_{W}$ 5.22 $\sigma_{Zh}^{365}$ 6.24
$\mathcal{S}_{HB}$ 5.28 $m_{W}$ 4.11 $\sigma_{WW}^{240}$ 5.71
$\mathcal{S}_{H}$ 4.84 $m_{W}$ 2.80 $\sigma_{Zh}^{240}$ 4.97
$\mathcal{S}_{2G}$ 3.87 $m_{W}$ 3.49 $R_{\mu}^{163}$ 4.40
$\mathcal{S}_{B-W}$ 3.38 $A^{Z}_{e}$ 3.23 $\sigma_{Zh}^{365}$ 3.94
$\mathcal{S}_{6}$ 0.39 $m_{W}$ 0.99 $\sigma_{Zh}^{240}$ 0.99
@@
$\mathcal{S}_{g}$ 0.47 $m_{W}$ 0.00 0.47
Table 5: $2\sigma$ CL in the SILH basis. Coefficients are renormalized at $\Lambda=$1\text{\,}\mathrm{TeV}$\$